An air-gap between two pole pieces is 20 mm in length and the area of the flux path across the gap is 5 cm2. If the flux required in the air-gap is 0.75 mWb. Find the mmf necessary.
1
Expert's answer
2020-06-18T14:51:47-0400

Given

An air-gap between two pole pieces is 20 mm in length ,

A = 5 cm2 which is 5*10-4 m2

Air gap then = 0.75 mwb = 0.75*10-3

Solution

we know that B = uoH

m.m.f = H*L

but H = (alpha/Auo) * L and Uo = 4*3.14*10-7

=(0.75*10-3* 20*10-3)/(5*10-4 * 4 *3.14 * 10-7)

= 0.238*105 A

Therefore m.m.f = 23870 A